OpenStudy (anonymous):

smarties help ;* I need a few ways to write the equations for slope. Im making a study guide and this is recommend to study but I'm not really sure what they are. Please help.

OpenStudy (anonymous):

what are the slops

OpenStudy (anonymous):

what?

OpenStudy (curiousquestions):

I can help. Just give me a sec.

OpenStudy (anonymous):

awesome thx @CuriousQuestions

OpenStudy (curiousquestions):

np

OpenStudy (curiousquestions):

So, you know how to find the slope, right?

OpenStudy (anonymous):

i just need to know the equations not how to solve them

OpenStudy (curiousquestions):

Is this what you need? \[Slope=\frac{ y^2-y^1 }{x^2- x^1 }\]

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

i need three ways to write the equations of slope

OpenStudy (curiousquestions):

Ok.

OpenStudy (curiousquestions):

Btw, instead of using the word `Slope` use `m`.

OpenStudy (curiousquestions):

\[y-y_1=m(x-x_1)\] & \[y=mx\]

OpenStudy (igreen):

I eat smarties, I don't talk to them.

OpenStudy (anonymous):

ha me too but if your smart i need helpp

OpenStudy (conqueror):

I'll help you devour your smarties if that's what you want.

OpenStudy (igreen):

If you are looking for equations for slope, here are some: \(\sf y = mx + b\) Slope-intercept form Where 'm' is the slope and 'b' is the y-intercept. \(\sf y - y_1 = m(x - x_1)\) Where 'm' is the slope, and (x1, y1) is a point on the line.

OpenStudy (xmissalycatx):

One way is Slope Intercept Formula where the slope is m y=mx + b Another way is by finding two points on a graph (x1 and y1 and x2 and y2) or \[slope = \frac{ \Delta x }{ \Delta y }\]

OpenStudy (igreen):

If you want to find slope, you can use the equaation: \(\sf m = \dfrac{y_2-y_1}{x_2-x_1}\) Where (x1, y1) and (x2, y2) are points given.

OpenStudy (xmissalycatx):

@iGreen is 3 smart 5 me :')

OpenStudy (anonymous):

Kalayn321 i like you alot

OpenStudy (anonymous):

Thnaks @iGreen

OpenStudy (igreen):

Yw.